OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is. Can someone help me with this?

OpenStudy (anonymous):

to find F, u can distribute the 5/9 into the parenthesis for part 2 plug in 20 for the eqn u get for part 1

OpenStudy (radar):

C = 5/9 (F-32) Part 1. C=5F/9-(5/9)32 5F/9=C+160/9 F=(9/5)C+(9/5)(160/9) F=(9/5)C+32

OpenStudy (radar):

Part 2. As instructed by amk11 substitute 20 for C F=(9/5)20+32 F=36+32=68 degrees Fahrenheit

OpenStudy (radar):

For fun find out F when it is -40 degrees Celsius!